OpenStudy (anonymous):

Which answer shows this equation in standard form? plzz help 8 – 4(x – y) = –6x + 6 A. –10x + 4y = –2 B. 2x + 4y = –2 C. 2x + 4y = 2 D. 2x – 4y = –2

OpenStudy (anonymous):

Well standard form looks like this:\[Ax+By=C\] So we just need to rearrange the expression so that it is in that order. We'll start with the original expression: \[8 – 4(x – y) = –6x + 6\]Subtract the 6 from both sides\[ – 4(x – y) = –6x\]Then we distribute the 4\[2 – 4x +4y = –6x\]add -6x to both sides\[2+2x +4y = 0\]and now we subtract the two from both sides\[2x +4y = -2\]So the answer is B

OpenStudy (anonymous):

wat are the question marks

OpenStudy (anonymous):

thx so much do you want a medal and ill fan you

OpenStudy (anonymous):

thx for the help

OpenStudy (anonymous):

No problem!